OpenStudy (timaashorty):

Help Please ! What is the number of real solutions ? 12x^2-11x=11 A. Two real solutions B. one real solution C. cannot be determined D. no real solutions

OpenStudy (shamil98):

Tima, the apply the quadratic formula.

OpenStudy (shamil98):

12x^2 -11x - 11 = 0 x = -b +- √b^2 -4ac

OpenStudy (shamil98):

plug the values in and find the solutions

OpenStudy (luigi0210):

Use the power within! But first check the discriminant :)

OpenStudy (shamil98):

oh yeah that works too b^2 - 4ac

OpenStudy (timaashorty):

Guys just give me the answer, im timed!

OpenStudy (timaashorty):

Then explain after words. lol

OpenStudy (shamil98):

-.-

OpenStudy (shamil98):

I'm no longer glazing your donut! it's two real solutions.

OpenStudy (luigi0210):

That urge to report..

OpenStudy (timaashorty):

Hehe Thanks Shamil. LOL And Luigi Im not gunna give a medal cuz of that.-.-

OpenStudy (shamil98):

Now, you are going to stay and read my explanation. okay?

OpenStudy (shamil98):

b^2 - 4ac If discriminant > 0 quadratic equation will have two real roots. If discriminant = 0 quadratic equation will have two equal roots. If discriminant < 0 quadratic equation will have no real roots (imaginary)

OpenStudy (shamil98):

so in this case (-11)^2 - 4(12)(-11) = 121 + 528 = 649 the discriminant is positive.

OpenStudy (luigi0210):

Pfft, who needs medals?

OpenStudy (shamil98):

so it has two real roots.

OpenStudy (shamil98):

UNDERSTAND TIMA?

OpenStudy (timaashorty):

LATER SHAMIL. Ill Read everything you wrote I rpomise.

OpenStudy (timaashorty):

HEY IT LET ME GIVE YOU BOTH MEDALS. Whaa??

OpenStudy (timaashorty):

Engagement: New Title! Learner Thats What I Got !

OpenStudy (shamil98):

*reports*

OpenStudy (luigi0210):

\[\Huge \mathbb{\color{green}{CONGRATULATIONS}} \color{red}{\ddot\smile}\]

OpenStudy (timaashorty):

^^^^^^^^ THANK YOU! (:

OpenStudy (timaashorty):

*Undo's Shamils medal*